Henrys Law constant for most gases decreases with increased temperature. Given that KH for O2 is 2.20 x 10-3 M atm-1 at 5oC, calculate the percent O2(aq) saturation if a lake water sample at 5oC and sea-level is found to contain, 9.20 mg/L of O2(aq). (Recall: partial pressure of O2 at sea-level = 0.2095 atm).
1
Expert's answer
2019-12-09T08:04:33-0500

Saturated concentration = "K_Hx"


"2.2\times 10^{-3}\times0.2095\\=4.609\times10^{-4}M"

9.2 mg per litre = "\frac{9.2}{32}\times10^{-3}= 2.875\times10^{-4}M"

Percent saturation= "\frac{2.875}{4.609}\times100=62.37\%"